Solve the following system of equations using a matrix.
3x-5y+z=-8
4x+y-z=46
5x-7y+3z=0

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Solve the following system of equations using a matrix.
3x-5y+z=-8
4x+y-z=46
5x-7y+3z=0
----
Use any method you know to get:
x = 10
y = 8
z = 2
==========
Cheers,
Stan H.
==========

Answer by Alan3354(69443)

It's not solved using a matrix, it is a matrix.